Materials used by the Instrument Division of Ziegler Inc. are currently purchased from outside suppliers at a cost of $1,350 per unit. However, the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Instrument Division at a variable cost of $900 per unit.

Assume that a transfer price of $1,200 has been established and that 75,000 units of materials are transferred, with no reduction in the Components Division's current sales.

a. How much would Ziegler Inc.'s total operating income increase?
$

b. How much would the Instrument Division's operating income increase?
$

c. How much would the Components Division's operating income increase?
$
Answer #1

Answer a

Ziegler Inc.'s total operating income increased by $ 33,750,000

Explanation:

Net benefit = (1350-900)*75,000

= 33,750,000

Answer b

Instrument Division's operating income increased by $ 11,250,000.

Explanation:

Net benefit = (1350-1200)*75,000

= 11,250,000

Answer c

Components Division's operating income increased by $ 22,500,000.

Explanation:

Net benefit = (1200-900)*75,000

= 22,500,000

In case of any doubt, please comment.